convert to decimal: 1). 4 3/5 2). 2/3 3).75.3%

Found 2 solutions by radh, kevwill:
Answer by radh(108)

1. Every 1 is equal to 100%, so it'd be 460%. (3/5=6/10=60/100)

2. 66.7% (Rounded; the 6 would go on forever otherwise)

3. 753/1000 (Because of the decimal, the percentage is no longer out of a simple 100, but is out of 1000).

Hope this helps!


Answer by kevwill(135)

I think the original solver got this one wrong. The problem is to convert each of the above numbers to a decimal, but they converted the first two to percentages and the third to a fraction.
The correct answers are:
1) 4.6
2) 0.666... (the 6 repeats forever, so it could be shown as 0.6 with a bar over the 6)
3) 0.753